OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 5x. What is the value of (f o g)(3)?

OpenStudy (anonymous):

f(g(3)) = f(5(3)) = f(15) = -2(15)+8 = -22

OpenStudy (anonymous):

thank you :)